OpenStudy (anonymous):

1 = 2ln(900/900-x) Solve for x

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (kc_kennylau):

Divide both sides by 2

OpenStudy (anonymous):

ok

OpenStudy (kc_kennylau):

man do it

OpenStudy (anonymous):

so 1/2 = ln(900/900-x) raise each side to e

OpenStudy (anonymous):

e^(1/2) = 900/(900-x)

OpenStudy (anonymous):

900-x = 900/e^(1/2)

OpenStudy (anonymous):

x = 900 - 900/e^91/2)

OpenStudy (anonymous):

?

OpenStudy (kc_kennylau):

Congratulations :D

OpenStudy (anonymous):

answer in my book is Po = 900(1-e^(-6))

OpenStudy (kc_kennylau):

Can you give me the original question?

OpenStudy (anonymous):

Time of extinction = 2ln(900=/900 - x) I used x instead of Po sorry. If time of extinction = 1 year find x (Po which is initial population.)

OpenStudy (kc_kennylau):

hmm...

OpenStudy (kc_kennylau):

maybe your book is wrong?